# Distinguishing Orbiting and Infalling Dark Matter Particles with Machine Learning

### v1

@@ Abstract
1. Introduction
In a $\Lambda$CDM Universe, halos are the key building blocks of structure within which galaxies form (Rees & Ostriker, 1977). Quantifying this general picture of matter in the Universe requires that we define some measure of what belongs to a halo. For example, halo masses allow us to summarize the abundance and growth of structure via mass functions (Press & Schechter, 1974) and underpin our understanding of the galaxy-halo connection (Wechsler & Tinker, 2018).
@@ 1. Introduction
The conventionally favored solution has been to define halos with spherical overdensity (SO) boundaries (Lacey & Cole, 1994), meaning radii chosen to enclose some specific overdensity, e.g. an overdensity of $200$ times the mean or critical density of the Universe (denoted $R_{\rm 200m}$ and $R_{\rm 200c}$, respectively). While measuring any definition of halo mass is observationally challenging, SO definitions can at least be easily measured in simulations, namely from spherically averaged density profiles. However, SO definitions suffer from a number of serious drawbacks. First, they do not include the outer regions of all orbits, leading to erroneously classified “backsplash” subhalos (and satellites) outside the alleged halo boundary (e.g., Balogh et al., 2000; Mamon et al., 2004; Knebe et al., 2020; Diemer, 2021) which contributes to the phenomenon of assembly bias (Gao et al., 2005; Villarreal et al., 2017; Mansfield & Kravtsov, 2020). Second, SO radii and masses partially evolve via pseudo-evolution due to the evolving background density of the Universe (Cuesta et al., 2008; Diemer et al., 2013). And third, a sharp halo boundary leads to an unphysical break in halo models of the large-scale distribution of matter (e.g., Cooray & Sheth, 2002; García et al., 2021).
@@
These issues have motivated a search for alternative, more physical definitions. One option is the splashback radius, which is defined to include the orbits of the most recently accreted particles and subhalos, avoiding the issue of backsplash halos and pseudo-evolution (Diemer & Kravtsov, 2014; Adhikari et al., 2014; More et al., 2015). However, due to the distribution of orbital properties, splashback radii are not uniquely defined either, and measuring them can be hard for halos with few particles. More fundamentally, the splashback radius still represents a “hard” boundary, which leads to some undesirable effects. Besides subtle questions of subhalo assignments (e.g., Garcia & Rozo, 2019), halos can be far from spherical, even out to large radii (e.g., Jing & Suto, 2002; Mansfield et al., 2017).
@@
While gravitational boundness is often used to define the matter that constitutes a halo, it is ill-defined in a Universe without a uniquely defined potential at infinity (Behroozi et al., 2013c; Voit, 2025). One commonly used boundary-free definition is the set of particles within a certain “linking length” of one another, an algorithm called friends-of-friends (or FOF, Davis et al., 1985). While simple computationally, there is no clear physical motivation for FOF halos, especially since the linking length is an arbitrary parameter and does not correspond to a particular SO definition (More et al., 2011).
@@
Based on particle dynamics, we can devise a more natural, boundary-free definition that is physically motivated even for non-spherical halos: a halo is the collection of particles orbiting within its potential well. A particle’s first pericentric passage serves as a physically motivated criterion to distinguish when it leaves the infalling population and begins orbiting the halo. This definition has recently been explored as a guiding principle for understanding halos and their density profiles (Diemer, 2022, 2023, 2025; Garcia et al., 2022; Salazar et al., 2024). While the orbits of dark matter particles are not observable, the satellite population can at least approximately be understood as a mixture of orbiting and infalling populations (Adhikari et al., 2021; Aung et al., 2021). However, some pressing questions remain, for example how to define subhalos in the orbiting-infalling framework or whether orbiting-only halos exhibit a more universal mass function.
@@
Such investigations are necessarily based on a classification of particles as orbiting or infalling, often for billions of particles. A simple cut in $r-v_{r}$ space (radius and radial velocity, respectively) is insufficient because a negative radial velocity does not provide enough information to decide whether a particle is on its first infall or has already experienced an apocenter. Information such as the time when particles fell into a halo helps but does not completely separate the populations (Garcia et al., 2022). Thus, Diemer (2022) presented an algorithm that considers the full historical information of each particle’s orbit. While reliable, this algorithm is computationally intensive and cannot easily be applied to the largest simulations, where often only a subset of particles or snapshots can be stored (e.g., Garrison et al., 2018).
@@
The paper is organized as follows. In Section 2, we describe our simulations, our training dataset, and the training of our model. In Section 3, we analyze the performance of the model in various regions of the input parameter space as well as with density profiles. We further interpret the physics behind the model’s decisions in Section 4 using feature importance and compare the model to other approximate methods using cuts in kinetic-energy space. We summarize our findings in Section 5.
2. Methods
2.1. Simulations
We use six $N$-body simulations from the Erebos suite (Diemer & Kravtsov, 2014, 2015), which model the same $\Lambda$CDM cosmology, with the WMAP7 parameters, that was used in the Bolshoi simulation ($\Omega_{\rm m}=0.27$, $\Omega_{\rm b}=0.0469$, $\sigma_{8}=0.82$, $n_{\rm s}=0.95$, Komatsu et al., 2011). The simulation box sizes range from $63\>{\rm Mpc}$ (L0063) to $2000\>{\rm Mpc}$ (L2000), each simulated with $1024^{3}$ particles. Finally, we use the L0125, L0250, and L0500, $1024^{3}$ particle simulations with a Planck-like cosmology ($\Omega_{\rm m}=0.32$, $\Omega_{\rm b}=0.0491$, $\sigma_{8}=0.834$, and $n_{\rm s}=0.9624$, Planck Collaboration et al., 2014) to assess how well the model generalizes to other cosmologies. For all simulations, we use the phase-space halo finder Rockstar (Behroozi et al., 2013a) to extract halo information from each simulation, with halo trees constructed using the Consistent-Trees code (Behroozi et al., 2013b).
We use the orbiting-infalling classifications based on the algorithm of Diemer (2022), which is implemented in the Sparta framework (Diemer, 2017, 2020), as our fiducial definition. Sparta labels a particle as orbiting if it has undergone one pericenter and as infalling otherwise. A particle is deemed to have had a pericenter when its radial velocity switches from negative to positive. To avoid spurious pericenters caused by noise, an additional constraint is imposed: the first passage requires the angular direction, defined as $\phi(t)=\hat{\textbf{r}}(t)\cdot\hat{\textbf{r}}(t_{ini})$, to be less than 0.5 at the snapshot before or after the particle switches in radial velocity. The majority of pericenters are determined via these conditions, although the algorithm also handles some special cases (for details, see Diemer, 2022). As a refinement of the original algorithm, we enforce an additional maximum physical velocity of $1.5\ V_{\rm 200m}$ at the particle’s pericenter, where $V_{\rm 200m}=\sqrt{GM_{\rm 200m}/R_{\rm 200m}}$ and $M_{\rm 200m}$ is the mass of the halo enclosed within $R_{\rm 200m}$. This condition prevents rare particles with excessively high velocities from being assigned pericenters, as they are unlikely to remain bound to the halo. We do still see orbiting particles with larger velocities in our datasets as this condition is only applied at the particle’s first pericenter.
@@ 2.1. Simulations
2.2. Training Dataset
We consider all particles within $4R_{\rm 200m}$ of each halo for our dataset, as Diemer et al. (2017) found that this radius contains the vast majority of orbiting particles. We limit our analysis to host halos with a minimum of 200 particles, as consistently defining when subhalo particles have a pericenter around their host halo is challenging (e.g., Han et al., 2012).
@@ 2.2. Training Dataset
We calculate the particles’ physical velocities and positions relative to the center of the host halo to use as model inputs. Due to the isotropy of the halos we can fully define a particle’s physical velocity in terms of its radial and tangential components. The physical velocity of each particle is $\vec{v}=\vec{v}_{\rm ptl}-\vec{\rm v}_{\rm halo}+\vec{v}_{\rm hub}$ (where $\vec{v}_{\rm hub}=H_{0}r$ is the Hubble velocity), the radial velocity is $v_{\rm r}=\vec{v}\cdot\hat{r}$, and the tangential velocity is $v_{\rm t}=\sqrt{v^{2}-v_{\rm r}^{2}}$. We then normalize the radii by $R_{\mathrm{200m}}$ and velocities by $V_{\mathrm{200m}}$. The radius, radial velocity, and tangential velocity from an initial snapshot are the first three input parameters for our model.
@@
The remaining inputs are the same three parameters ($r/R_{\rm 200m}$, $v_{\rm r}/V_{\rm 200m}$, and $v_{\rm t}/V_{\rm 200m}$) evaluated at a past snapshot. The addition of the past trajectory information allows the model to distinguish between particles that converged to the same phase-space region at the current snapshot despite traveling distinct paths. To maximize the the range of halos present in the simulations we chose our primary snapshot to be at $z\sim 0$. We have tested time separations between the primary and past snapshot of 0.5, 0.75, 1, and 2 dynamical times (where $t_{\rm dyn}=2R_{\rm 200m}/V_{\rm 200m}$). We find that the model performs best with the snapshots separated by 0.75 dynamical times, with the present snapshot at $z=0.031$ and the past snapshot at $z=0.395$. We tested models using no past snapshots which led to noticeable decline in performance and multiple past snapshots which provided to negligible improvement, indicating that two snapshots are sufficient to accurately classify orbiting and infalling particles. A side effect of using multiple snapshots is that some particles are not within our search radius both at the present and the past snapshot, or the halo they belong to in the current snapshot did not exist in the past. These particles form a significant population, and we include them in our dataset with an indication that they are “missing” values for their past parameters. The model still learns from these particles (see Section 4.1).
2.3. Model Training
@@
For our machine learning model, we select XGBoost (Chen & Guestrin, 2016) because it is easily implemented, scalable to large datasets, and interpretable. XGBoost makes classifications with an ensemble of gradient-boosted decision trees. Each decision tree consists of a series of binary splits based on the particle’s input values (e.g., $r/R_{\rm 200m}\leq 0.5$) until it reaches the final node, which assigns a probability of whether the particle is orbiting or infalling. The final ensemble decision is a combination of each of the tree’s probabilities, weighted by the performance of the tree during training. Since each decision tree consists of a series of simple splits, the model’s decisions can be easily analyzed in depth, providing insights into the particle populations that constitute the halo.
@@
To train the decision trees, XGBoost uses gradient boosting, constructing each tree to minimize the errors from the prior tree. XGBoost provides user settings that control overfitting and training time, called the learning rate and subsample rate. We use the default learning rate of 0.3, which limits the impact of each new tree’s prediction on the ensemble classification. Instead of the default subsample rate of 1, we use a rate of 0.5, so each decision tree only receives 50% of the dataset. This results in shorter training times and reduces the risk of overfitting, as each decision tree is trained on a random subset of data. During training, XGBoost is able to learn from our missing feature values, such as the absence of past particle properties in prior snapshots, incorporating this information into its classifications (Section 2.2).
@@ 3. Results
In this section, we evaluate the accuracy of our model’s classifications by comparing them against classifications made by the Sparta-based particle trajectory method in the validation halo sample (Section 2.2). We compare the predictions as a function of phase-space parameters (Section 3.1) and in stacked density profiles (Section 3.2).
3.1. Accuracy Across Phase Space
@@
Figure 4.— Median total (left), orbiting (center), and infalling (right) density profiles of all particles from the WMAP7 test dataset. We bin halos by peak height and hide radial bins where fewer than 50% of the total number of halos contain particles, with darker colors representing larger peak heights. The solid lines are density profiles from the model’s classifications, and the dashed lines are the profiles from Sparta. The bottom panels show the corresponding ratio between the median profiles, with the shaded region representing the 16th and 84th percentile of the ratios between the individual halo profiles. Since both our total profiles and SPARTA’s profiles are constructed from the raw particle count in each radial bin, they must agree, as confirmed in the left panel. The model accurately picks up on the key characteristics of the orbiting and infalling profiles across masses. The increased errors for larger peak heights are largely results of a smaller population of halos.
@@
There are outlier orbiting particles with extreme velocities reaching ±$\pm 10v_{\rm r}/V_{\rm 200m}$ and $10v_{\rm t}/V_{\rm 200m}$ (D9–D12). A large portion of the orbiting particles at large radii with positive radial velocities are “fly-bys”, i.e., particles that have had a pericenter but do not remain bound to the halo. The remaining particles are truly orbiting and eventually return to the halo. These outlier regions contain few particles and thus do not significantly impact the training of the model nor our evaluation of its performance.
At the past snapshot (right column of Fig. 2), the orbiting and infalling particles are distributed similarly to the present snapshot, albeit with three key differences. First, not all particles have a second snapshot (Section 2.2), reducing the overall phase-space density. Second, particles in the past snapshot have a broader range of radial velocities at smaller radii. This phenomenon reflects the broad range of paths that particles can take to arrive at the present snapshot’s configuration. Third, distant infalling particles with smaller radial velocities ($r/R_{\rm 200m}>1.5$ and $-0.5<v_{\rm r}/V_{\rm 200m}<0.25$ in panel D8) typically cannot reach the halo in time to experience a pericenter. This manifests as a low-density region in panel D12, splitting the orbiting particles into two distinct tails at large radii.
@@
The second population consists of infalling particles outside the center of the halo ($10.5<R_{\rm 200m}<$ 1) with outgoing radial velocities, manifesting as an arc in panel S5. These radii and radial velocities are also characteristic of orbiting particles approaching their apocenters, contributing to increased misclassification in this region. We suspect that subhalo interactions deflect a small fraction of infalling particles into this region. However, the low occupation of infalling particles in this region mitigates the impact of the high misclassification rates on the overall performance.
3.2. Accuracy of Density Profiles
One of the primary goals of the separation of particles into orbiting and infalling was to understand the physical nature of the respective density profiles (Diemer, 2022), making density profiles a crucial benchmark for our particle classification method. In this section, we compare the model’s density profiles with Sparta’s for the WMAP7 simulations on which the model was trained. We then evaluate the model’s ability to generalize to other cosmologies.
We compare the density profiles from our model’s classifications with the profiles from Sparta in Fig. 4. To identify potential mass biases in the model’s predictions, we further split the profiles by peak height, $\nu_{\Delta}$, which reflects the statistical significance of halos in the linear overdensity field. We define $\nu_{\Delta}=\delta_{c}(z)/\sigma_{M_{\Delta}}$, where $\delta_{c}(z)=1.686$ is the overdensity threshold used by the spherical top-hat collapse model (Gunn & Gott, 1972) and $\sigma_{M_{\Delta}}$ is the variance of the linear power spectrum measured within spheres of the Lagrangian radius of the halo. The orbiting component dominates the overall profile at small radii, gradually decreasing until a sharp cutoff. On the other hand, the infalling profile’s contribution is negligible near the center and overtakes the orbiting contribution at large radii. For a detailed analysis of the orbiting and infalling density profiles, refer to Diemer (2022) and Salazar et al. (2024).

### v2

@@ I. Introduction
The conventionally favored solution has been to define halos with spherical overdensity (SO) boundaries (lacey_94), meaning radii chosen to enclose some specific overdensity, e.g. an overdensity of $200$ times the mean or critical density of the Universe (denoted $R_{\rm 200m}$ and $R_{\rm 200c}$, respectively). While measuring any definition of halo mass is observationally challenging, SO definitions can at least be easily measured in simulations, namely from spherically averaged density profiles. However, SO definitions suffer from a number of serious drawbacks. First, they do not include the outer regions of all orbits, leading to erroneously classified “backsplash” subhalos (and satellites) outside the alleged halo boundary (e.g., balogh_00; mamon_04; knebe_20; diemer_21_subs) which contributes to the phenomenon of assembly bias (gao_05_ab; villareal_17; mansfield_20_ab). Second, SO radii and masses partially evolve via pseudo-evolution due to the evolving background density of the Universe (cuesta_08; diemer_13_pe). And third, a sharp halo boundary leads to an nonphysical break in halo models of the large-scale distribution of matter (e.g., cooray_02; garcia_21).
@@
These issues have motivated a search for alternative, more physical definitions. One option is the splashback radius, which is defined to include the orbits of the most recently accreted particles and subhalos, avoiding the issue of backsplash halos and pseudo-evolution (diemer_14; adhikari_14; More et al., 2015). However, due to the distribution of orbital properties, splashback radii are not uniquely defined either, and measuring them can be hard for halos with few particles. More fundamentally, the splashback radius still represents a “hard” boundary, which leads to some undesirable effects. Besides subtle questions of subhalo assignments (e.g., Garcia and Rozo, 2019), halos can be far from spherical, even out to large radii (e.g., jing_02; mansfield_17).
@@
While gravitational boundness is often used to define the matter that constitutes a halo, it is ill-defined in a Universe without a uniquely defined potential at infinity (behroozi_13_unbound; voit_25). One commonly used boundary-free definition is the set of particles within a certain “linking length” of one another, an algorithm called friends-of-friends (or FOF, davis_85). While simple computationally, there is no clear physical motivation for FOF halos, especially since the linking length is an arbitrary parameter and does not correspond to a particular SO definition (more_11_fof).
@@
Based on particle dynamics, we can devise a more natural, boundary-free definition that is physically motivated even for non-spherical halos: a halo is the collection of particles orbiting within its potential well. A particle’s first pericentric passage serves as a physically motivated criterion to distinguish when it leaves the infalling population and begins orbiting the halo. This definition has recently been explored as a guiding principle for understanding halos and their density profiles (diemer_22_prof1; diemer_23_prof2; diemer_25_prof3; Garcia et al., 2022; Salazar et al., 2024). While the orbits of dark matter particles are not observable, the satellite population can at least approximately be understood as a mixture of orbiting and infalling populations (adhikari_21; Aung et al., 2021). However, some pressing questions remain, for example how to define subhalos in the orbiting-infalling framework or whether orbiting-only halos exhibit a more universal mass function.
@@
Such investigations are necessarily based on a classification of particles as orbiting or infalling, often for billions of particles. A simple cut in $r-v_{r}$ space (radius and radial velocity, respectively) is insufficient because a negative radial velocity does not provide enough information to decide whether a particle is on its first infall or has already experienced an apocenter. Information such as the time when particles fell into a halo helps but does not completely separate the populations (Garcia et al., 2022). Thus, diemer_22_prof1 presented an algorithm that considers the full historical information of each particle’s orbit. While reliable, this algorithm is computationally intensive and cannot easily be applied to the largest simulations, where often only a subset of particles or snapshots can be stored (e.g., garrison_18).
@@
The paper is organized as follows. In Section II, we describe our simulations, our training dataset, and the training of our model. In Section III, we analyze the performance of the model in various regions of the input parameter space as well as with density profiles. We further interpret the physics behind the model’s decisions in Section IV using feature importance and compare the model to other approximate methods using cuts in kinetic-energy space. We summarize our findings in Section V.
II. Methods
II.1. Simulations
[…]
We use the orbiting-infalling classifications based on the algorithm of diemer_22_prof1, which is implemented in the Sparta framework (Diemer, 2017, 2020), as our fiducial definition. Sparta labels a particle as orbiting if it has undergone one pericenter and as infalling otherwise. A particle is deemed to have had a pericenter when its radial velocity switches from negative to positive. To avoid spurious pericenters caused by noise, an additional constraint is imposed: the first passage requires the angular direction, defined as $\phi(t)=\hat{\textbf{r}}(t)\cdot\hat{\textbf{r}}(t_{ini})$, to be less than 0.5 at the snapshot before or after the particle switches in radial velocity. The majority of pericenters are determined via these conditions, although the algorithm also handles some special cases (for details, see diemer_22_prof1). As a refinement of the original algorithm, we enforce an additional maximum physical velocity of $1.5\ V_{\rm 200m}$ at the particle’s pericenter, where $V_{\rm 200m}=\sqrt{GM_{\rm 200m}/R_{\rm 200m}}$ and $M_{\rm 200m}$ is the mass of the halo enclosed within $R_{\rm 200m}$. This condition prevents rare particles with excessively high velocities from being assigned pericenters, as they are unlikely to remain bound to the halo. We do still see orbiting particles with larger velocities in our datasets as this condition is only applied at the particle’s first pericenter.
@@ II.1. Simulations
II.2. Training Dataset
We consider all particles within $4R_{\rm 200m}$ of each halo for our dataset, as Diemer et al. (2017) found that this radius contains the vast majority of orbiting particles. We limit our analysis to host halos with a minimum of 200 particles. We exclude subhalos because it is challenging to consistently determine if their constituent particles have a pericenter (e.g., Han et al., 2012), which is required by our definition of orbiting.
@@ II.2. Training Dataset
We calculate the particles’ physical velocities and positions relative to the center of the host halo to use as model inputs. Due to the average isotropy of the halos we can capture most of the information about a particle’s physical velocity in terms of its radial and tangential components. The physical velocity of each particle is $\vec{v}=\vec{v}_{\rm ptl}-\vec{\rm v}_{\rm halo}+\vec{v}_{\rm hub}$ (where $\vec{v}_{\rm hub}=H(z)\times r$ is the Hubble velocity), the radial velocity is $v_{\rm r}=\vec{v}\cdot\hat{r}$, and the tangential velocity is $v_{\rm t}=\sqrt{v^{2}-v_{\rm r}^{2}}$. With the halo’s $R_{\mathrm{200m}}$ and $V_{\mathrm{200m}}$ at the current snapshot, we normalize the particles’ radii and velocities respectively. The radius, radial velocity, and tangential velocity from an initial snapshot are the first three input parameters for our model.
@@
The remaining inputs are the same three parameters ($r/R_{\rm 200m}$, $v_{\rm r}/V_{\rm 200m}$, and $v_{\rm t}/V_{\rm 200m}$) evaluated at a past snapshot. The addition of the past trajectory information allows the model to distinguish between particles that converged to the same phase-space region at the current snapshot despite traveling distinct paths. To maximize the the range of halos present in the simulations we chose our primary snapshot to be at $z\sim 0$. We have tested time separations between the primary and past snapshot of 0.5, 0.75, 1, and 2 dynamical times (where $t_{\rm dyn}=2R_{\rm 200m}/V_{\rm 200m}$). We find that the model performs best with the snapshots separated by 1.0 dynamical time, with the present snapshot at $z=0.031$ and the past snapshot at $z=0.567$. We tested models using no past snapshots which led to noticeable decline in performance and multiple past snapshots which provided negligible improvement, indicating that two snapshots are sufficient to accurately classify orbiting and infalling particles. A side effect of using multiple snapshots is that some particles are not within our search radius both at the present and the past snapshot, or the halo they belong to in the current snapshot did not exist in the past. These particles form a significant population, and we include them in our dataset with an indication that they are “missing” values for their past parameters. The model still learns from these particles (see Section IV.1).
II.3. Model Training
@@
For our machine learning model, we select XGBoost (Chen and Guestrin, 2016) because it is easily implemented, scalable to large datasets, and interpretable. XGBoost makes classifications with an ensemble of gradient-boosted decision trees. Each decision tree consists of a series of binary splits based on the particle’s input values (e.g., $r/R_{\rm 200m}\leq 0.5$) until it reaches the final node, which assigns a probability of whether the particle is orbiting or infalling. The final ensemble decision is a combination of each of the tree’s probabilities, weighted by the performance of the tree during training. Since each decision tree consists of a series of simple splits, the model’s decisions can be easily analyzed in depth, providing insights into the particle populations that constitute the halo.
@@
To train the decision trees, XGBoost uses gradient boosting, constructing each tree to minimize the errors from the prior tree. XGBoost provides user settings that control overfitting and training time, called the learning rate and subsample rate. We use the default learning rate of 0.3, which limits the impact of each new tree’s prediction on the ensemble classification. Instead of the default subsample rate of 1, we use a rate of 0.5, so each decision tree only receives 50% of the dataset. This results in shorter training times and reduces the risk of overfitting, as each decision tree is trained on a random subset of data. During training, XGBoost is able to learn from our missing feature values, such as the absence of past particle properties in prior snapshots, incorporating this information into its classifications (Section II.2).
@@ III. Results
In this section, we evaluate the accuracy of our model’s classifications by comparing them against classifications made by the Sparta-based particle trajectory method in the validation halo sample (Section II.2). We compare the predictions as a function of phase-space parameters (Section III.1) and in stacked density profiles (Section III.2).
III.1. Accuracy Across Phase Space
@@
Figure 4.— Median orbiting (left), and infalling (right) density profiles of all particles from the WMAP7 test dataset. We bin halos by peak height and hide radial bins where fewer than 50% of the total number of halos contain particles, with darker colors representing larger peak heights. The solid lines are density profiles from the model’s classifications, and the dashed lines are the profiles from Sparta. The bottom panels show the corresponding ratio between the median profiles, with the shaded region representing the 16th and 84th percentile of the ratios between the individual halo profiles. Since both our total profiles and SPARTA’s profiles are constructed from the raw particle count in each radial bin, they must agree, as confirmed in the left panel. The model accurately picks up on the key characteristics of the orbiting and infalling profiles across masses. The increased errors for larger peak heights are largely results of a smaller population of halos.
@@
There are outlier orbiting particles with extreme velocities reaching $\pm 10v_{\rm r}/V_{\rm 200m}$ and $10v_{\rm t}/V_{\rm 200m}$ (D9–D12). A large portion of the orbiting particles at large radii with positive radial velocities are “fly-bys”, i.e., particles that have had a pericenter but do not remain bound to the halo. The remaining particles are truly orbiting and eventually return to the halo. These outlier regions contain few particles and thus do not significantly impact the training of the model nor our evaluation of its performance.
At the past snapshot (right column of Fig. 2), the orbiting and infalling particles are distributed similarly to the present snapshot, albeit with three key differences. First, not all particles have a second snapshot (Section II.2), reducing the overall phase-space density. Second, particles in the past snapshot have a broader range of radial velocities at smaller radii. This phenomenon reflects the broad range of paths that particles can take to arrive at the present snapshot’s configuration. Third, distant infalling particles with smaller radial velocities ($r/R_{\rm 200m}>1.5$ and $0.5<v_{\rm r}/V_{\rm 200m}<0.25$ in panel D8) typically cannot reach the halo in time to experience a pericenter. This manifests as a low-density region in panel D12, splitting the orbiting particles into two distinct tails at large radii.
@@
The second population consists of infalling particles outside the center of the halo ($0.5<R_{\rm 200m}<1$) with outgoing radial velocities, manifesting as an arc in panel S5. These radii and radial velocities are also characteristic of orbiting particles approaching their apocenters, contributing to increased misclassification in this region. We suspect that subhalo interactions deflect a small fraction of infalling particles into this region. However, the low occupation of infalling particles in this region mitigates the impact of the high misclassification rates on the overall performance.
III.2. Accuracy of Density Profiles
One of the primary goals of the separation of particles into orbiting and infalling was to understand the physical nature of the respective density profiles (diemer_22_prof1), making density profiles a crucial benchmark for our particle classification method. In this section, we compare the model’s density profiles with Sparta’s for the WMAP7 simulations on which the model was trained. We then evaluate the model’s ability to generalize to other cosmologies.
We compare the density profiles from our model’s classifications with the profiles from Sparta in Fig. 4. To identify potential mass biases in the model’s predictions, we further split the profiles by peak height, $\nu_{\Delta}$, which reflects the statistical significance of halos in the linear overdensity field. We define $\nu_{\Delta}=\delta_{c}/\sigma_{M}(M,z)$, where $\delta_{c}=1.686$ is the overdensity threshold used by the spherical top-hat collapse model (gunn_72) and $\sigma_{M}(M,z)$ is the variance of the linear power spectrum measured within spheres of the Lagrangian radius of the halo. The orbiting component dominates the overall profile at small radii, gradually decreasing until a sharp cutoff. On the other hand, the infalling profile’s contribution is negligible near the center and overtakes the orbiting contribution at large radii. For a detailed analysis of the orbiting and infalling density profiles, refer to diemer_22_prof1 and Salazar et al. (2024).